OpenStudy (anonymous):

Can anyone help? So confused!! a hexagonal tray of vegetables has an area of 450 cm^2. What is the length of each side of the hexagon? Also, what is the area at the bottom in terms of the side length?

OpenStudy (anonymous):

Does anyone know how to do this??

OpenStudy (anonymous):

@phi can you help?

OpenStudy (phi):

do you know the formula for the area of a hexagon ?

OpenStudy (anonymous):

no I don't

OpenStudy (phi):

see wikipedia http://en.wikipedia.org/wiki/Hexagon#Regular_hexagon for a formula that gives the area as a function of the length of a side

OpenStudy (anonymous):

ok

OpenStudy (phi):

do you see the formula ?

OpenStudy (anonymous):

yes, but it sure looks confusing.

OpenStudy (phi):

\[ A = \frac{3\sqrt{3}}{2} t^2 \] where t is the length of 1 side

OpenStudy (phi):

multiply both sides by the reciprocal of 3sqrt(3)/2 to get t^2 by itself

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so would I end up with 450 cm^2 times t^2

OpenStudy (anonymous):

oh wait, I messed up. nevermind

OpenStudy (phi):

write down your steps

OpenStudy (anonymous):

2/3^3(450cm^2) = 2/3^3(3^3/2)

OpenStudy (phi):

and don't forget the side call it s^2 (not sure why wiki uses t^2) 2/3^3(450cm^2) = s^2 (the other stuff cancels out) now simplify the left side

OpenStudy (phi):

\[ \frac{2 \cdot 450}{3 \sqrt{3} } = s^2 \]

OpenStudy (anonymous):

900/3^3 = s^2

OpenStudy (anonymous):

oops I didnt mean 3^3 I meant 3 sqrt 3

OpenStudy (phi):

you could divide 3 into 900, right ? and I would multiply top and bottom by sqrt(3) to move the sqrt up top

OpenStudy (anonymous):

yeah 3 into 900 is 300

OpenStudy (phi):

multiply top and bottom by sqrt(3)

OpenStudy (anonymous):

1558.8/9

OpenStudy (anonymous):

900 * sqrt 3 = 1558.8 and 3 sqrt 3 * sqrt 3 = 9

OpenStudy (phi):

I would not change the answer to decimals, as that is not an exact answer you should get \[s^2= 100 \sqrt{3} \]

OpenStudy (anonymous):

ohh ok

OpenStudy (phi):

now take the square root of both sides

OpenStudy (phi):

btw, \[ \sqrt{\sqrt{x}} = \sqrt[4]{x} \]

OpenStudy (anonymous):

I think I just confused myself. lol so I would take the square root of 100 sqrt 3?

OpenStudy (phi):

and s^2 on the other side

OpenStudy (anonymous):

10^4√3 ?

OpenStudy (phi):

yes, and now you can give the decimal approximation \[ s= 10 \sqrt[4]{3} \approx 13.161\]

OpenStudy (anonymous):

yay!

OpenStudy (phi):

btw, could also write it as \[s= 10 \cdot 3^{\frac{1}{4}} \] that is 3 to the 1/4 power

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

Thanks for your help! That made understand it alot better.

OpenStudy (anonymous):

made me^

OpenStudy (phi):

If you could google for the formula, you could have figured it out. divide the hexagon into 6 congruent triangles, and figure out the area of one of the triangles. then multiply by 6 to get the hexagon

OpenStudy (anonymous):

oh ok. google just looked real confusing when I first saw the formula, but that makes more sense now.